2 1/6 - 1 1/4 and reduce to lowest form
Answer #1

2 1/6 = (2*6+1)/6 = 13/6

1 1/4 = (4*1+1)/4= 5/4

13/6 - 5/4 = 13*2 - 5*3 / 12

= 11/12